OpenStudy (anonymous):

is it possible to find acceleration with only coefficient of kinetic friction and it's mass -.-

OpenStudy (amistre64):

hmm, well, since Force = mass x accel and Force_k = uk x n im thinking maybe

OpenStudy (amistre64):

n = mass*gravity*cos(theta) Force_k = mass x (uk*gravity*cos(theta) .... maybe

OpenStudy (amistre64):

a more descriptive question might prove better

OpenStudy (amistre64):

\[\mu_k=tan\theta-\frac ag sec\theta\] \[\frac ag sec\theta=tan\theta-\mu_k\] \[a=g~cos\theta~(tan\theta-\mu_k)\] \[a=g~(sin\theta~-\mu_kcos\theta)\]